a garden plot is to certain 200 sq. ft. if its length is to be twice its width. what should its dimension be?
Answer by FrankM(1040)

Width = X
Length = 2X
L*W = area, so 2X*X = 2X^2 = 200
X^2 = 100
X=10
2X=20